OpenStudy (anonymous):

Identify the solution of this inequality x+4/ x-8 is greater than/ equal to 0

OpenStudy (ankit042):

x>8 U x<=-4

OpenStudy (jdoe0001):

how would you solve for "x" say => \(\cfrac{x+4}{x-8} = 0\)

OpenStudy (anonymous):

x=-4 and x=8

OpenStudy (anonymous):

our answer choices are 8 or none of the above

OpenStudy (jdoe0001):

hmm not sure on the x = 8, if you cross multiply, the x-8 goes poof and all you're left with is x+4 which gives you x =-4 well, on an inequality is more or less the same so the inequality will be \(\bf x \ge -4\)

OpenStudy (anonymous):

How about this one. 4x+3 / x-5. What do we do with the 4x+3?

OpenStudy (jdoe0001):

so let's try using 8 x = 8 \(\bf 8 \ge -4\) so that's true and is valid HOWEVER remember the original equation was => \(\bf \cfrac{x+4}{x-8} = 0\) if we use x = 8 that makes the denominator 8-8 = 0 and a fraction with a denominator of 0 is undefined, so we have that restriction

OpenStudy (jdoe0001):

so x has to be greater than 8, or less than 8, but can't be 8, because we end up with an undefined expression now if we use x = -4, then look at the numerator x + 4, x = -4 => -4 + 4 = 0 0/WHATEVER = 0, and that's equal to 0, so that's ok if you go below -4, you'd get a negative number in the numerator and positive in the denominator that yields a negative fraction, which is not greater than or equal to 0

OpenStudy (jdoe0001):

so based on the restrictions, "x" has to be greater than -4 but not 8 \(\bf x>-4; \ \ x \ne 8\)

OpenStudy (ankit042):

I think X > -4 will not work...try putting x =-3 and check your solution

OpenStudy (jdoe0001):

\(\bf \cfrac{-3+4}{-3-8} \implies \cfrac{1}{-11} \lt 0\)

OpenStudy (jdoe0001):

true

OpenStudy (jdoe0001):

I gather then that x has to be greater than or equal to 0, but not 8

OpenStudy (ankit042):

The correct solution for this problem is X <=-4 Union x>8 Let me know if anyone needs a detailed solution

OpenStudy (jdoe0001):

\(\bf \bf x \ge 0; \ \ x \ne 8\)

OpenStudy (jdoe0001):

hmmm, now that I think about it, the I have a -8 in the denominator :( yes you're right @ankit042

OpenStudy (jdoe0001):

so there @sherifoster , as @ankit042 originally said, yeah, based on the restrictions it has to be that \(\bf x \le -4 \cup x > 8\)

OpenStudy (jdoe0001):

if you use something like -5 -7 .... the numerator becomes negative, so does the denominator and thus the fraction will be positive and greater than 0 if you use a value greater than 8, both numerator and denominator's are positive and that 's greater than 0 any value in between those 2 intervals, will give use either undefined at 8, or a negative atop or bottom, any negative rational is of not greater or equal to 0